OpenStudy (anonymous):

: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)(2x^2y^6)

OpenStudy (anonymous):

multiply all the constant terms and add the powers of like terms

OpenStudy (anonymous):

multiplying similar bases result in summin the power of these similar bases (3x^3y^4)(2x^2y^6) = 3*2(x^(3+2) y^(4+6) ) = 6 x^5 y^10

OpenStudy (lalaly):

\[6x^{3+2}y^{4+6} = 6x^5y^{10}\]